OpenStudy (anonymous):

what is the simplified form of... (x+4/x^2-3x-10) • (x-3/x^2+x-12)

sam (.sam.):

In this problem 2*-5=-10 and 2-5=-3, so insert 2 as the right hand term of one factor and -5 as the right-hand term of the other factor. ((x+4)/((x+2)(x-5)))((x-3)/(x^(2)+x-12)) 4*-3=-12 and 4-3=1, so insert 4 as the right hand term of one factor and -3 as the right-hand term of the other factor. ((x+4)/((x+2)(x-5)))((x-3)/((x+4)(x-3))) Reduce the expression by canceling out the common factor of (x-3) from the numerator and denominator. ((x+4)/((x+2)(x-5)))(1)/(x+4) Multiply ((x+4))/((x+2)(x-5)) by (1)/((x+4)) 1/(x+2)(x-5)